OpenStudy (anonymous):

Three-eighths of the teachers at Naran's school are men, and 30 of the teachers are women. (Each teacher is either a man or a woman.) How many teachers are there at Naran's school?

OpenStudy (anonymous):

If 3/8 of teachers are men, then 5/8 are women. If 30 = 5/8, work out how many people would be 3/8.

OpenStudy (anonymous):

Let t be the total number of teachers.\[\frac{5 t}{8}=30,\text{ }t=48 \]

OpenStudy (anonymous):

Thankyou for the medal.

OpenStudy (anonymous):

your welcome!